OpenStudy (anonymous):

Thanks for helping! 6. Prove the identity: (cosx+cosy)^2+(sinx-siny)^2=2+2cos(x+y) Part I: Complete the left-hand column of the table below following the steps indicated in the right-hand column. (12 points) Given on the left side of the original problem (cosx+cosy)^2+(sinx-siny)^2 Expand each squared term Simplify the expression Apply the Pythagorean identity Apply the addition formula for cosine

OpenStudy (anonymous):

@dan815 @Kainui @jim_thompson5910 @zepdrix

OpenStudy (dan815):

expand left side

OpenStudy (dan815):

(cosx+cosy)^2+(sinx-siny)^2=??

OpenStudy (anonymous):

2+2cos(x+y)

OpenStudy (dan815):

no i mean expand it

OpenStudy (anonymous):

oh gotcha

OpenStudy (dan815):

and use the cos trig identities

OpenStudy (dan815):

cosxcosy+sinxsiny = cos(x+y) ?

OpenStudy (anonymous):

cos^2(x)+2cos(x)cos(y)-2sin(x)sin(x)+cos^2(x)+sin^2(y)+sin^2

OpenStudy (anonymous):

That is the identity?

OpenStudy (dan815):

i dunno maybe its a negative inbetween

OpenStudy (dan815):

cosxcosy- sinxsiny = cos(x+y) ?

OpenStudy (dan815):

http://prntscr.com/7daa8n

OpenStudy (dan815):

check it out

OpenStudy (anonymous):

thanks, from here I can simplify and use the Pythagorean identity?

OpenStudy (dan815):

yep

OpenStudy (dan815):

sin^2(a) + cos^2(a) = 1;

OpenStudy (dan815):